OpenStudy (anonymous):

please help? For a particular sample of 58 scores on a psychology exam, the following results were obtained. First quartile = 52 Third quartile = 84 Standard deviation = 10 Range = 66 Mean = 79 Median = 78 Mode = 81 Midrange = 65 Answer each of the following: I. What score was earned by more students than any other score? Why? II. What was the highest score earned on the exam? III. What was the lowest score earned on the exam? IV. According to Chebyshev's Theorem, how many students scored between 59 and 99? V. Assume that the distribution is normal. Bas

OpenStudy (cruffo):

I. Mode = # occurring most often (could be more than one number)

OpenStudy (anonymous):

is there any way to tell for sure?

OpenStudy (cruffo):

For I ?? It is the definition of "mode".

OpenStudy (cruffo):

http://www.purplemath.com/modules/meanmode.htm

OpenStudy (anonymous):

ohh ok. That makes sense. I just wasn't sure because for the mode there was no way of telling which number occured more often.

OpenStudy (cruffo):

the problem tells you which score occurred most often: Mode = 81

OpenStudy (anonymous):

ohh ok. I really got tripped up and started to over analyze. Thank you

OpenStudy (anonymous):

for the highest answers, do we use the mean?

OpenStudy (cruffo):

not sure, I'm thinking about it...

OpenStudy (anonymous):

ok. Thank you

OpenStudy (cruffo):

highest score is 98.

OpenStudy (cruffo):

I used the mid-range and the range to find it.

OpenStudy (anonymous):

what do you mean? Adding them together?

OpenStudy (cruffo):

Let H = highest score Let L = lowest score Midrange = \(\large \dfrac{H +L}{2}\) and Range = \(\large H-L\)

OpenStudy (anonymous):

ohh! ok!

OpenStudy (cruffo):

we know that \(H-L = 66\) so \( L = H - 66\). We also know that \( \dfrac{H +L}{2} = 65\) so substituting \(H - 66\) for \(L\) we get \[ \frac{H+H-66}{2} = 65\] \[\frac{2H - 66}{2} = 65\] \[2H-66 = 130\] \[2H = 196\] \[H = 98\]

OpenStudy (anonymous):

2H would be 2*H, right?

OpenStudy (cruffo):

Right, 2H means 2 times H.

OpenStudy (anonymous):

that is awesome. Thank you tons for helping me. It's very tricky

OpenStudy (cruffo):

yah, I wasn't sure how to get that at first...

OpenStudy (anonymous):

you're a lifesaver :D

OpenStudy (cruffo):

lol :)

OpenStudy (sirm3d):

how about that last question?